OpenStudy (anonymous):

my question is: 2x^+3x-2=3

OpenStudy (anonymous):

2x^2 + 3x -2 -3 = 0 2x^2 + 3x -5 = 0 2x^2 + 5x - 2 x - 5 x(2x+5) -(2x+5) (x-1) (2x+5)

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

np! Good luck! Ask for any clarrifications.

OpenStudy (anonymous):

hello um i dont understand how it became 5x when it was originally just -5 ?

OpenStudy (anonymous):

Hi! Your breaking up the middle term. 2x^2 + 3x -2 -3 = 0 Now what i just did was add -2 and -3 = -5. 5x comes from decomposition of the middle term 3x